# Repeatable Energy-Efficient Perching for Flapping-Wing Robots Using Soft Grippers

### v1

@@ 3.3 Flight Testing
Following the static and free-fall tests, a series of free-flight tests were conducted to demonstrate the system’s ability to perform a full perch cycle in a realistic environment. The generic perching cycle can hereby be divided into three phases, as clarified by Fig. 11). This cycle is initiated from the flying condition, with the landing phase, where the system approaches the perching object. This is followed, after successful attachment, by a perched phase, which enables the system to perform a defined mission or enter a state of rest. Lastly, the take-off phase is performed from a standstill, and the flying condition is again entered, completing the cycle.
@@
During testing, our system’s first phase (green strip in Fig. 11) begins in a generic free-flying condition. It consists of landing the drone on the perch through deceleration and a nearly vertical approach path, so that the gripping mechanism triggers upon contact and attaches the drone securely to the branch. In the second phase (blue strip), the servo-motor is automatically disconnected from power (500 ms after impact) to preserve energy, and the drone is passively held atop the branch, with the motors shut off. This phase can be extended for as long as needed, depending on the mission. In the third phase (yellow strip), the drone’s motors are rapidly engaged, and the take-off command is given in the form of an IR signal. This signal resets the ratchet mechanism and simultaneously opens the grippers to make take-off possible. Take-off is then typically performed near-vertically; however, a degree of deviation of up to 36° to the vertical is permitted, corresponding to the shape of the opened grippers.
@@
The full perching cycle could be successfully performed several times, in a continuous take. Video data was recorded both at 500 fps, in slow motion, and at 60 fps at 4K, with simultaneous IMU data recording for all tests. Video footage of these flight tests can be found in Extension 1.3 for an exemplary slow-motion recording and in Extension 1.4 for the real-time recordings of the full perching cycle.
@@
An example of a successful continuous perch-cycle can be seen in Fig. 11. The IMU Z-acceleration data is plotted against time-from-impact for the approach phase, and against time-from-release for the take-off phase. For compactness, the data for the perched state, where acceleration remains constant, is partly cut out. However, the data shown in the plot comes from a single recording, and the system stayed perched in a stable manner without power to any motors until the take-off sequence was initiated. In Extension 1.5 the full acceleration data of this uninterrupted perching cycle can be seen. To reduce the noise and obtain clearer results, the IMU data was filtered through a lowpass Butterworth filter. Two filters were considered, one with a cut-off frequency of 20 Hz, still allowing for the flapping frequency of 19 Hz to be observed, and the other with a lower cut-off frequency of 7 Hz, thus excluding the flapping effects and presenting a clearer view.
@@
A more detailed view of the collected data from a free-flight sequence during landing and subsequent take-off is shown in Fig. 12 and Fig. 13, respectively.
@@
Fig. 12.B shows the filtered Z-acceleration data over time, starting with the initiation of the landing sequence. Here, a large impact acceleration can be seen upon contact, where a more sophisticated trigger setup could be implemented to reduce this peak. However, the current design was able to withstand these forces without failure. Fig. 12.A displays the integrated acceleration data, thus showing the velocity during approach and impact, all the way until standstill. Here, the steady approach and the deceleration time frame can be seen.
@@
The take-off sequence in Fig. 13 begins at release time. As the flight trajectory is diagonal in the X-Z-plane (see coordinates in Fig 11), the acceleration data is plotted in X and Z components (panel B). Panel A shows the integrated combined acceleration, i.e. the velocity during take-off until a continuous flight speed of 1.48 m/s is reached. Due to the quick opening of the grippers upon receiving the release signal, the take-off sequence can be carried out with any desired start velocity, provided the flapping-wing MAV is given sufficient thrust to take flight.
@@
The free-flight tests showed that the system enables a flapping-wing MAV of over 170 g to land on a perch, attach to it in a stable manner, remain attached for an extended period of time, while conserving energy, and reliably take off when desired, thus completing a full perch cycle.
@@
During flight testing, a few strategies for achieving a successful perching attempt and subsequent take-off could be noted. In the approach, it was beneficial to position the MAV in a zone atop the perching object with the gripper axis and perching object axis in parallel. Then, in a powered descent, the alignment of these axes in a vertical plane was conducted, and upon achieving near coincidence, the throttle could be reduced to zero, letting the system enter a free-fall until contact with the perch. However, it should be noted that this approach strategy was helpful for manual steering and could be optimised and parallelised in an automated approach. Still, approximate axis alignment as a priority followed by a rapid descent would also be suitable for a simple automated system. For take-off, the timing of gripper release is of primary importance; here, the best strategy is to quickly engage the throttle, to reduce disturbances to the stability introduced by a slow flapping start, and upon achieving at least enough power to enter a hover state, the grippers can be safely opened again. In an automated system, this could be easily implemented and optimised timing-wise for a quick take-off.
@@
While the emphasis of the current study was on demonstrating the full flight cycle, and thus testing was conducted by a pilot, the manoeuvre could potentially be automated, as mentioned, making the framework more functional and robust. Doing so would first require adapting the control system to handle the higher payload and the resulting change in flight behaviour. An analysis of the effects of an increased payload on the flight stability and performance of lightweight flapping-wing robots with perching capabilities can be found in van Bruegge and Armanini, (2025).
@@ 4 Conclusion
In this paper, we presented a framework for energy-efficient perching in mid-size tailless FWMAVs based on a bioinspired soft-gripping mechanism. The viability and performance of the framework were extensively tested in over two hundred perching tests of the integrated system, in manual tests, free-fall tests and controlled free flight. Furthermore, the complete functionality and full perch cycle of landing, continuous perching and take-off could be repeatedly demonstrated and recorded. The gripping mechanism itself weighs below 39 g and was successfully integrated into an exemplary tailless FWMAV platform, the Flapper Nimble+, resulting in a system with roughly 170 g take-off mass. The lightweight and soft grippers used for perching attach themselves to different types of objects without damaging the surface, while a custom trigger system ensures reliable and fast activation of the gripping process upon contact. A precise ratchet mechanism allows the robot to passively maintain a stable position on the perch without a power drain and enables take-off from the perch thanks to an integrated reset mechanism. With this framework, the full perch cycle of landing, perching and take-off can be repeated at will.
Future work will focus on developing control approaches to ensure safe and stable flight of lightweight flapping-wing robots with perching capabilities despite the considerable added payload of a gripping mechanism, as well as on achieving fully autonomous take-off and perched landing. Combining advanced control strategies with the proposed perching framework will significantly expand the versatility and capabilities of flapping-wing robots, taking them yet again a step closer to real-world applicability.
Appendix A Appendix
A.1 Index to Multimedia Extensions

### v2

@@ 3.3 Flight Testing
Following the static and free-fall tests, a series of free-flight tests were conducted to demonstrate the system’s ability to perform a full perch cycle in a realistic environment. The generic perching cycle can hereby be divided into three phases, as clarified by figure 14. This cycle is initiated from the flying condition, with the landing phase, where the system approaches the perching object. This is followed, after successful attachment, by a perched phase, which enables the system to perform a defined mission or enter a state of rest. Lastly, the take-off phase is performed from a standstill, and the flying condition is again entered, completing the cycle.
@@
During testing, our system’s first phase (green strip in figure 14) begins in a generic free-flying condition. It consists of landing the drone on the perch through deceleration and a nearly vertical approach path, so that the gripping mechanism triggers upon contact and attaches the drone securely to the branch. In the second phase (blue strip), the servo-motor is automatically disconnected from power (500 ms after impact) to preserve energy, and the drone is passively held atop the branch, with the motors shut off. This phase can be extended for as long as needed, depending on the mission. In the third phase (yellow strip), the drone’s motors are rapidly engaged, and the take-off command is given in the form of an IR signal. This signal resets the ratchet mechanism and simultaneously opens the grippers to make take-off possible. Take-off is then typically performed near-vertically; however, a degree of deviation of up to 35° to the vertical is permitted, corresponding to the shape of the opened grippers.
@@
The perching cycle was successfully performed several times in separate tests for landing and take-off and has also been demonstrated in continuous takes to show the viability of the system. Video data was recorded both at 500 fps, in slow motion, and at 60 fps at 4K, with simultaneous IMU data recording for all tests. Video footage of these flight tests can be found in Extension 1.3 for an exemplary slow-motion recording and in Extension 1.4 for the real-time recordings of the full perching cycle.
@@
An example of a successful continuous perch-cycle can be seen in figure 14. The IMU Z-acceleration data is plotted against time-from-impact for the approach phase, and against time-from-release for the take-off phase. For compactness, the data for the perched state, where acceleration remains constant, is partly cut out. However, the data shown in the plot comes from a single recording, and the system stayed perched in a stable manner without power to any motors until the take-off sequence was initiated. In Extension 1.5, the full acceleration data of this uninterrupted perching cycle can be seen. To reduce the noise and obtain clearer results, the IMU data was filtered through a low-pass Butterworth filter. Two filters were considered, one with a cut-off frequency of 20 Hz, still allowing for the flapping frequency of 19 Hz to be observed, and the other with a lower cut-off frequency of 7 Hz, thus excluding the flapping effects and presenting a clearer view.
@@
A more detailed view of the collected data from a free-flight sequence during landing and subsequent take-off is shown in figure 15 and figure 16, respectively.
@@
Figure 15.B shows the filtered Z-acceleration data over time, starting with the initiation of the landing sequence. Here, a large impact acceleration can be seen upon contact, where a more sophisticated trigger setup could be implemented to reduce this peak. However, the current design was able to withstand these forces without failure. Figure 15.A displays the integrated acceleration data, thus showing the velocity during approach and impact, all the way until standstill. Here, the steady approach and the deceleration time frame can be seen.
@@
The take-off sequence in figure 16 begins at release time. As the flight trajectory is diagonal in the X-Z-plane (see coordinates in figure 14), the acceleration data is plotted in X and Z components in panel B. Panel A shows the integrated combined acceleration, i.e. the velocity during take-off until a continuous flight speed is reached. Due to the quick opening of the grippers upon receiving the release signal, the take-off sequence can be carried out with any desired start velocity, provided the flapping-wing MAV is given sufficient thrust to take flight.
@@
The free-flight tests showed that the system enables a flapping-wing MAV of over 170 g to land on a perch, attach to it in a stable manner, remain attached for an extended period of time, while conserving energy, and reliably take off when desired, thus completing a full perch cycle. While attaching a gripper to the FWMAV introduces a penalty on the flight time, the capability to ”rest” on a perch or conduct a perched mission adds value and extends the flight time. As the goal of our first prototype was to demonstrate the viability of the perching operation, a significant reduction in flight time was accepted. However, this is mostly due to the external control board and the unoptimised gripper prototype, which can be made lighter in future work.
@@
During flight testing, a few strategies for achieving a successful perching attempt and subsequent take-off could be noted. In the approach, it was beneficial to position the MAV in a zone atop the perching object with the gripper axis and perching object axis in parallel. Then, in a powered descent, the alignment of these axes in a vertical plane was conducted, and upon achieving near coincidence, the throttle could be reduced to zero, letting the system enter a free-fall until contact with the perch. However, it should be noted that this approach strategy was helpful for manual steering and could be optimised and parallelised in an automated approach. Still, approximate axis alignment as a priority, followed by a rapid descent, would also be suitable for a simple automated system. For take-off, the timing of gripper release is of primary importance; here, the best strategy is to quickly engage the throttle, to reduce disturbances to the stability introduced by a slow flapping start, and upon achieving at least enough power to enter a hover state, the grippers can be safely opened again. In an automated system, this could be easily implemented and optimised timing-wise for a quick take-off.
@@
While the emphasis of the current study was on demonstrating the full flight cycle, and thus testing was conducted by a pilot, the manoeuvre could potentially be automated, as mentioned, making the framework more functional and robust. Doing so would first require adapting the control system to handle the higher payload and the resulting change in flight behaviour. An analysis of the effects of an increased payload on the flight stability and performance of lightweight flapping-wing robots with perching capabilities can be found in van Bruegge and Armanini (2025).
@@ 4 Conclusion
In this paper, we presented a framework for energy-efficient perching in mid-size tailless FWMAVs based on a bioinspired soft-gripping mechanism. The viability and performance of the framework were extensively tested in over two hundred perching tests of the integrated system, in manual tests, free-fall tests and controlled free flight. Furthermore, the complete functionality and full perch cycle of landing, continuous perching and take-off could be repeatedly demonstrated and recorded. The gripping mechanism itself weighs below 39 g and was successfully integrated into an exemplary tailless FWMAV platform, the Flapper Nimble+, resulting in a system with roughly 170 g take-off mass. The lightweight and soft grippers used for perching attach themselves to different types of objects without damaging the surface, while a custom trigger system ensures reliable and fast activation of the gripping process upon contact. A precise ratchet mechanism allows the robot to passively maintain a stable position on the perch without a power drain and enables take-off from the perch thanks to an integrated reset mechanism. With this framework, the full perch cycle of landing, perching and take-off can be repeated at will. Future work will focus on developing control approaches to ensure safe and stable flight of lightweight flapping-wing robots with perching capabilities despite the considerable added payload of a gripping mechanism, as well as on achieving fully autonomous take-off and perched landing. Combining advanced control strategies with the proposed perching framework will significantly expand the versatility and capabilities of flapping-wing robots, taking them yet again a step closer to real-world applicability.
Appendix
.1 Index to Multimedia Extensions
@@
Acknowledgements
The authors received no financial support for the research, authorship, and/or publication of this article.